# Portable, Massively Parallel Implementation of a Material Point Method for Compressible Flows

### v2

@@ 5 Discussion and conclusion
This paper has showcased the potential of a massively parallel, portable implementation of the Material Point Method for compressible flows. Insight has been given on algorithm adaptations which will provide further benefits, moreover, the simulation of standard test cases of compressible flow has provided results comparable with the literature and with established software runs with the same polynomial order (first order shape functions - Lagrangian hat functions) and number of degrees of freedom. These results show that MPM is able to simulate compressible gas dynamics, giving qualitatively good results with a prototype developed modifying existing implementations [65, 14], and serve as further step towards the design of a highly efficient MPM-based transonic and supersonic FSI solver.
@@
Performance tests on NVIDIA A100 GPUs have confirmed that the main computational steps of the MPM algorithm are the Particles-to-Grid and Grid-to-Particles kernels, and that they scale well with the size of the problem. Moreover, adding a reordering step is beneficial to data locality and, consequently, performance, while data races in the P2G kernel can be tackled CUDA’s atomics at acceptable cost, even though alternative exist. Last, this approach has shown to be easily portable to different hardware architectures, as different vendors GPUs and CPUs.
@@
  • On the modelling side, more complex test cases could be simulated, eventually leading to a 3D extension and Fluid-Structure Interaction models. In addition, the effect on parallel performance of higher order B-Splines under development will need to be carefully investigated.
    A 3D extension will facilitate a further growth in computational load, but this will not introduce conceptual main changes in the code, also thanks to the high level, portable approach, which does not refer explicitly to specific, small size memories. However, it could justify a more efficient I/O, e.g., moving from JSON to HDF5. Simulations characterised by a higher computational load would also further benefit from a future multi-GPUs approach.
@@
    In particular, stencil growth due to expansion of shape function support to neighbouring grid elements will introduce halo overlapping, leading to an increase in communication costs, as well as to a constraint on the maximum level of achievable parallel granularity.
@@
  • For the algorithm implementation, the use of atomic operations to handle data races was chosen for simplicity, also because it showed an acceptable impact on computational time and scalability on NVIDIA A100 GPUs and a simple portability to other architectures (section 4). To reach the full parallelisation capabilities proved by the G2P kernels also for the P2G ones, P2G algorithm adaptations will be necessary. An attractive and and more portable route would involve avoiding atomics, and writing a grid colouring algorithm that can be implemented in an architecture-independent way, relying only on standard libraries functions. Its main steps are:
@@
    In this way we would iterate on the vector holding the list of particles in each cell, and then inside each cell we would perform only a small amount of serial operations, without any data race.
  • On the High Performance Computing side, avenues of development include the support of multi-GPU and multi-node architectures via MPI, and a grid-particles adaptive refinement through a quad-tree approach [53, 15]. Moreover, currently the code employs Structure-of-Arrays for the memory, since, differently from arrays of arbitrary structures, it provides coalescing, which improves memory efficiency [30]. However, recent developments on GPUs computing literature [61, 22] show that the best performance is achieved by adopting an Array of SoA (AoSoA). When addressing the MPI-based multi-GPU implementation, this possibility could be explored: indeed, the current data structure gives a conceptually straightforward extension towards multi-GPU implementations, where a further level of parallelisation through an MPI-based domain decomposition will make each GPU hold a SoA as the current one, globally giving rise to an AoSoA data structure.

### v4

@@ 5 Discussion and conclusion
This paper has showcased the potential of a massively parallel, portable implementation of the Material Point Method for compressible flows. Insight has been given on algorithm adaptations which will provide further benefits, moreover, the simulation of standard test cases of compressible flow has provided results comparable with the literature and with established software runs with the same polynomial order ($\mathbb{Q}_{1}$ FEM basis functions on Cartesian tensor product grids) and number of degrees of freedom. These results show that MPM is able to simulate compressible gas dynamics, giving good results with a prototype which has been developed stemming from existing implementations [70, 13], and serve as further step towards the design of a highly efficient MPM-based transonic and supersonic FSI solver.
@@
Performance tests on NVIDIA A100 GPUs have confirmed that the main computational steps of the MPM algorithm are the Particles-to-Grid and Grid-to-Particles kernels, and that they scale well with the size of the problem. Moreover, adding a reordering step is beneficial to data locality and, consequently, to the performance, while data races in the P2G kernels can be tackled CUDA’s atomics at acceptable cost. Finally, it has been shown that an alternative to atomics exists, and that it could be explored in the look for a higher degree of performance portability, especially on CPU architectures. Ultimately, this approach has shown to be easily portable to different hardware architectures, as different vendors GPUs and CPUs.
@@
  • On the modelling side, more complex test cases could be simulated, eventually leading to a 3D extension and Fluid-Structure Interaction models. In addition, the effect on parallel performance of higher order B-Splines, currently under development, will need to be carefully investigated.
    A 3D extension will facilitate a further growth in computational load, but this will not introduce conceptual main changes in the code, thanks also to the high level, portable approach, which does not refer explicitly to specific, small size memories. Indeed, extension to the 3D case would require only slight modifications of the main algorithm, which would consist in repeating section 3 steps also in the z direction where necessary and in adding the relative variables to the mpm struct and the z component to the P2GD and G2PD kernels. In particular, the computational load of the 2D 16 million particles cylinder simulation of section 4 is exemplary of the one of 3D realistic test cases. Considering the extra memory in the z-direction, our 2D simulation performance is representative of the one of a 3D simulation with at least 10 million particles; a resolution which was used for recent, realistic scenarios MPM simulations in the literature [25]. As a matter of fact, adding one spatial dimension would not have a large impact on the code, since the largest share of the computational time is taken by kernels which we implement as Lagrangian particles-index based ones.
    However, as data files become larger, a 3D extension could justify a more efficient I/O, e.g., moving from JSON to HDF5. Simulations characterised by a higher computational load would also further benefit from a future multi-GPUs approach, that, to be optimal, would require the introduction of a load balance mechanism among the GPUs [35].
@@
    In particular, expansion of shape function support to neighbouring grid elements will introduce halo overlapping, leading to an increase in communication costs, as well as to a constraint on the maximum level of achievable parallel granularity.
@@
  • For the algorithm implementation, the use of atomic operations to handle data races has been chosen for simplicity, also because it has shown an acceptable impact on computational time and scalability on NVIDIA A100 GPUs and a simple portability to other architectures (section 4). To reach the full parallelisation capabilities proved by the G2P kernels also for the P2G ones, P2G algorithm adaptations will be necessary. An interesting and and more portable route would involve avoiding atomics, and writing a grid colouring algorithm that can be implemented in an architecture-independent way, relying only on standard libraries functions. Its main steps are:
@@
    This way we would iterate on the vector holding the list of particles in each cell, and then we would perform only a small amount of serial operations inside each cell, without any data race.
  • On the High Performance Computing side, avenues of development include the support of multi-GPU and multi-node architectures via MPI, and a grid-particles adaptive refinement through a quad-tree approach [59, 14]. Moreover, currently the code employs Structure-of-Arrays for the memory, since, differently from arrays of arbitrary structures, it provides coalescing, which improves memory efficiency [30]. However, recent developments on GPUs computing literature [66, 22] show that the best performance is achieved by adopting an Array of SoA (AoSoA). This possibility could be explored when addressing the MPI-based multi-GPU implementation. Indeed, the current data structure gives a conceptually straightforward extension towards multi-GPU implementations, where a further level of parallelisation through an MPI-based domain decomposition will make each GPU hold a SoA as the current one, globally giving rise to an AoSoA data structure.